### v3

# On Context Utilization in Summarization with Large Language Models

Abstract
Large language models (LLMs) excel in abstractive summarization tasks, delivering fluent and pertinent summaries. Recent advancements have extended their capabilities to handle long-input contexts, exceeding 100k tokens. However, in question answering, language models exhibit uneven utilization of their input context. They tend to favor the initial and final segments, resulting in a U-shaped performance pattern concerning where the answer is located within the input. This bias raises concerns, particularly in summarization where crucial content may be dispersed throughout the source document(s). Besides, in summarization, mapping facts from the source to the summary is not trivial as salient content is usually re-phrased. In this paper, we conduct the first comprehensive study on context utilization and position bias in summarization. Our analysis encompasses 5 LLMs, 10 datasets, and 5 evaluation metrics. We introduce a new evaluation benchmark called MiddleSum on the which we benchmark two alternative inference methods to alleviate position bias: hierarchical summarization and incremental summarization [^1] We will share all code and data upon publication..
1 Introduction
Large language models (LLMs) have drastically transformed the landscape of NLP recently Brown et al. (2020). With instruction tuning Ouyang et al. (2022); Chung et al. (2022), LLMs made a major leap forward in conditional (prompted) content generation, and can generate satisfying outputs without the need to finetune on a task specifically. In abstractive summarization specifically, this approach has arguably opened a new paradigm; summaries generated by LLMs are highly fluent, grammatical and relevant (Goyal et al., 2022). Despite noticeably lower scores on automatic metrics such as ROUGE (Lin, 2004) or BERTScore (Zhang et al., 2019), summaries generated by LLMs are largely preferred by humans over summaries from state-of-the-art fine-tuned models like BRIO (Liu et al., 2022b, 2023c). In fact, on XSum, GPT-3.5 summaries are even on par with re-annotated human-written summaries, and much better than the dataset ground-truth, according to human evaluators (Zhang et al., 2023b). LLMs also show promising capability in evaluating summaries generated by other systems, including other LLMs (Fu et al., 2023; Luo et al., 2023; Shen et al., 2023a).
Despite this success, a few major technological bottlenecks remain with LLMs, including the maximum length of their context window. The standard context window length for open-source LLMs is 2k tokens (Brown et al., 2020; Scao et al., 2022; Penedo et al., 2023; Touvron et al., 2023a), which drastically limits their usefulness for long-input summarization (Shaham et al., 2022). Several techniques were proposed to extend the context window, including ALiBi (Press et al., 2021), LeX (Sun et al., 2022), position interpolation (Chen et al., 2023) and YaRN (Peng et al., 2023). While some of them claim up to 128k tokens processing capacity (Peng et al., 2023), it remains unclear how much such methods help on long-context summarization.
Scaling up context length would only succeed if a key question gets addressed first: do LLMs make proper use of their entire context? Recent work (Liu et al., 2023a) suggested that, surprisingly, such a simple assumption may not hold: through experiments on multi-document question answering and key-value retrieval, the authors find that LLMs mostly focus on the beginning and end of the (long) context window. Plotting performance with regards to the position of the important information exhibits a U-shape, with performance high at first (beginning of the source), then dropping, and rising again at the end. Worryingly, in the middle of the context window, LLMs’ performance can drop to even below random chance, calling for greater examination of LLMs’ behaviors with regard to the position of information within the source.
In this work, we investigate in depth how LLMs use their context window in abstractive summarization. Unlike in question-answering, mapping facts in the output to a specific snippet in the source is not straightforward in abstractive summarization, due to the high-level of re-phrasing and compression. We conduct a large-scale study with 5 open-source LLMs, 10 datasets covering many aspects of summarization, and 5 highly diverse automatic metrics. Our contributions are threefold:
  • We conduct the first large-scale analysis on context utilization in abstractive summarization, and the impact of the position of salient information on performance. We show that the U-shape or middle-curse exhibited by (Liu et al., 2023a) also holds in abstractive summarization.
  • We craft an evaluation dataset (MiddleSum) where important information is concentrated in the middle of the context, enabling us to automatically quantify how much LLMs are affected by the middle-curse.
  • We benchmark two alternative methods for inference on MiddleSum: hierarchical summarization and incremental summarization, showing their promise at alleviating the middle curse (especially in the scientific paper domain).
2 Experimental Setup
Datasets
We cover a broad set of diverse abstractive summarization tasks, varying length and domain. We include 5 datasets of standard length (source is below 2k tokens, which always fits in the context window): (i) CNN/DailyMail (Hermann et al., 2015), (ii) XSum (Narayan et al., 2018), (iii) Reddit-TIFU (Kim et al., 2019), (iv) SAMSum (Gliwa et al., 2019), and (v) Multi-XScience (Lu et al., 2020). We also include another 5 long-input summarization datasets: (i) Arxiv and (ii) PubMed (Cohan et al., 2018), (iii) GovReport (Huang et al., 2021), (iv) SummScreenFD (Chen et al., 2022), and (v) Multi-News (Fabbri et al., 2019). A high-level view of each dataset is shown in Table 1, and detailed statistics are presented in Appendix A. For all datasets, we run experiments on the test set, subsampling 1,000 data points if its size is greater than 1,000, or using the entire test set otherwise.
Figure 1: Distribution of the relative location of summary bigrams within the source. We split each source document into 20 bins of the same number of words, and plot the distribution of summary bigrams over source bins.
Models
We experiment with several popular high-performing open-source instruction-tuned LLMs. Instruction-tuning datasets such as Flan (Wei et al., 2021) include some of the datasets we study: CNN/DM, XSum, SAMSum and Multi-News. We analyze the following models through HuggingFace transformers (Wolf et al., 2020):
  • Flan-UL2 is a 20B parameters encoder-decoder model pre-trained on 1T tokens. It is based on the UL2 20B model (Tay et al., 2022), with the addition of Flan-T5 (Chung et al., 2022) instruction fine-tuning. The context window is 2k.
  • Llama-2 (Touvron et al., 2023b) is a recently introduced powerful decoder-only model pre-trained on 2T tokens, ranging from 7B to 70B parameters, and with a 4k context window. We use the 7B and 13B models.
  • Xgen-7B (Nijkamp et al., 2023) is a 7B decoder-only model pre-trained on up to 1.5T tokens. It supports 8k tokens context window.
  • Mistral-7B (Jiang et al., 2023a) is also an 8k-context 7B decoder-only model, with performance slightly better than Llama-2-13B.
Input length # Documents In Flan?
Dataset Standard Long Single Multi Yes No
CNN/DM (Hermann et al., 2015)
XSum (Narayan et al., 2018)
Reddit-TIFU (Kim et al., 2019)
SAMSum (Gliwa et al., 2019)
Multi-XScience (Lu et al., 2020)
[…]
PubMed (Cohan et al., 2018)
GovReport (Huang et al., 2021)
SummScreenFD (Chen et al., 2022)
Multi-News (Fabbri et al., 2019)
Table 1: Summarization datasets under study. In standard length datasets, the context and summary fit within a 2k tokens LLM context window.
We use the instruction-tuned (or chat) checkpoints for Llama-2, Xgen-7B and Mistral-7B. To run inference, we use the following prompt: Read the following text and summarize it: [text]. Summarize the above text in [n] sentences. Summary: where n is set to an average number of target sentences per dataset (see Appendix A). We infer all models in bfloat16 and sample summaries with top-k sampling (Fan et al., 2018) using $k=50$.
Evaluation Measures
Summarization evaluation is especially challenging in the LLM era, as most automatic metrics poorly correlate with human preferences (Goyal et al., 2022; Liu et al., 2022a). To get a broad picture of performance, we evaluate with metrics as diverse as possible. First, we consider reference-based metrics: ROUGE-2 (Lin, 2004), which measures bigram overlap, BERTScore (Zhang et al., 2019), which measures semantic similarity with BERT (Devlin et al., 2018) embeddings, and A3CU (Liu et al., 2023c), which extracts facts in the form of Atomic Content Units (ACUs) Liu et al. (2022a), and checks the presence of ACUs between prediction and reference. As reference-free metrics, we include SummaC (Laban et al., 2022), a leading factual consistency evaluation metric relying on entailment scores between pairs of source and summary sentences. We also leverage GPT-3.5 (gpt-3.5-turbo-0125) as a summarization evaluator, which is proven to be a strong natural language generation evaluator (Wang et al., 2023a; Shen et al., 2023a; Jain et al., 2023). We prompt the model with the source and generated summary (which fits in GPT-3.5’s 16k context window) and ask to output a score on a likert scale from 1 to 5. We refer to Appendix B for the full prompt template. For GPT-3.5 evaluation, we subsample 300 data points per dataset in order to reduce costs.
We report the performance of LLMs on all 10 datasets, alongside a comparison to SOTA, in Appendix C. FLan-UL2 dominates on standard-length datasets, but Llama-2-13B has the upper hand on the long-input ones. Performance itself is not our focus in this paper, but rather which position-related factors influence it. We discard Flan-UL2 on long-input datasets due to poor performance.
Figure 2: Distribution of relative location of input context sentences aligned with sentences from summaries. X-axis corresponds to the source sentence bin, y-axis to the fraction of aligned sentences in each bin.
3 Experiments
In this section, we describe a series of experiments aimed at understanding how LLMs treat information in their input depending on the position.
3.1 RQ1: Where in the source do LLMs take their information from?
We investigate summaries generated by LLMs, and map them to specific parts of the input. Unlike in question-answering or extractive summarization, mapping salient information from a summary to the source is not trivial in abstractive summarization.
We follow the approach used in (Kim et al., 2019; Zhao et al., 2022a) and compute the relative position of bigams from generated summaries within the source documents, as a proxy for the position of salient information. We only use unique bigrams from summaries, and for each bigram, find all its occurrences within the source, if there are any. We then split the source into 20 bins of the same number of words, and compute the fraction of matched bigrams found in each bin. On top of the LLMs described above, we include the position of bigrams from reference summaries, and a uniform baseline.
As seen in Fig. 1, all summarization datasets except XSum, and Reddit-TIFU show some lead bias: salient bigrams from the reference (orange curves) are more likely to be found at the beginning of the source. However, LLMs show a significantly stronger lead bias on all datasets: bigrams from LLMs summaries are much more likely to be found in the first 20% words of the source. It is especially striking on XSum (except for Flan-UL2), Reddit-TIFU, Arxiv, PubMed and GovReport. On XSum, Flan-UL2 closely matches the reference distribution, which we attribute to its better instruction tuning. Results in Appendix D confirm that bigram distribution for LLMs and references are statistically different (p-value of Kolmogorov-Smirnov test (Massey Jr, 1951) inferior to 0.001) in all cases except Flan-UL2 on XSum and SAMSum and Llama-2-7B on XSum. We conclude that LLMs focus on contents at the beginning of the source document(s).
3.2 RQ2: Where do LLMs look at within their context window?
Metric Model
CNN/DM
XSum
Reddit
SAMSum
Multi-X
AVG
Arxiv
PubMed
GovReport
SummScreenFD
Multi-N
AVG
ROUGE-2 Flan-UL2 -0.296 -0.124 0.048 -0.069 -0.201 -0.128 _ _ _ _ _ _
Llama-2-7B -0.160 -0.023 0.063 -0.059 -0.100 -0.056 0.022 -0.113 -0.109 -0.079 -0.210 -0.098
Llama-2-13B -0.166 -0.086 0.031 -0.078 -0.039 -0.068 -0.017 -0.081 -0.166 -0.139 -0.213 -0.123
Xgen-7B -0.228 -0.042 0.066 -0.039 -0.041 -0.056 0.028 -0.091 -0.405 0.063 -0.283 -0.138
Mistral-7B -0.289 -0.031 0.006 -0.024 -0.052 -0.078 -0.270 -0.279 -0.585 -0.132 -0.324 -0.318
[…]
Llama-2-7B -0.173 -0.012 0.062 -0.130 -0.385 -0.128 -0.031 -0.203 -0.104 -0.067 -0.256 -0.132
Llama-2-13B -0.193 -0.102 0.038 -0.089 -0.352 -0.140 -0.082 -0.209 -0.063 -0.152 -0.279 -0.157
Xgen-7B -0.252 -0.106 0.046 -0.075 -0.343 -0.146 -0.017 -0.125 -0.353 -0.093 -0.345 -0.187
Mistral-7B -0.278 -0.052 0.014 -0.108 -0.416 -0.168 -0.348 -0.367 -0.567 -0.356 -0.403 -0.408
[…]
Llama-2-7B -0.182 -0.076 0.028 -0.121 -0.090 -0.088 -0.038 -0.209 -0.154 -0.129 -0.217 -0.149
Llama-2-13B -0.190 -0.098 0.009 -0.166 0.016 -0.086 -0.104 -0.232 -0.111 -0.198 -0.228 -0.175
Xgen-7B -0.212 -0.130 0.023 -0.126 -0.025 -0.094 -0.036 -0.255 -0.211 -0.076 -0.287 -0.173
Mistral-7B -0.291 -0.110 0.004 -0.105 -0.119 -0.126 -0.160 -0.283 -0.305 -0.010 -0.283 -0.208
SummaC Flan-UL2 -0.012 0.548 0.270 0.186 -0.035 0.191 _ _ _ _ _ _
Llama-2-7B 0.088 0.552 0.375 0.227 0.224 0.293 0.090 0.108 0.126 -0.020 0.205 0.102
Llama-2-13B 0.162 0.556 0.394 0.173 0.096 0.276 0.090 0.265 0.192 -0.144 0.232 0.127
Xgen-7B 0.001 0.161 0.220 0.117 0.004 0.101 -0.208 -0.087 -0.313 -0.141 0.046 -0.141
Mistral-7B -0.045 0.515 0.149 0.069 0.154 0.128 -0.250 -0.103 -0.387 0.124 -0.010 -0.125
[…]
Llama-2-7B 0.036 0.036 -0.152 0.077 -0.153 -0.031 0.008 -0.116 -0.013 -0.068 -0.120 -0.062
Llama-2-13B 0.039 0.072 -0.038 0.066 -0.052 0.017 -0.010 -0.084 -0.084 -0.060 -0.051 -0.058
Xgen-7B 0.056 -0.007 -0.101 0.006 -0.174 -0.044 -0.058 -0.096 -0.317 -0.063 -0.055 -0.118
Mistral-7B -0.115 0.124 -0.133 0.036 -0.204 -0.058 -0.446 -0.322 -0.580 -0.188 -0.163 -0.342
Table 2: Spearman correlation coefficient between each LLM’s metric, and the mean position of salient information within the context window. Flan-UL2 is not applied to long-context summarization datasets due to its too short context window. Multi-X is short for Multi-XScience, Multi-N is Multi-News dataset, AVG columns represent the average over standard-length and long-input datasets, respectively. Numbers in gray correspond to non-significant Spearman scores (p-value greater than 0.05).
In the previous experiment’s design, LLMs may not see the entire source in long-input summarization datasets, due to their limited context window, which is shorter than the source on the long-input datasets. We now focus on input information accessible to LLMs, and only consider salient information if it falls within the context window. Besides, since the same bigram may occur multiple times throughout the source, we adjust the methodology for saliency estimation. We align sentences in generated summaries to sentences in the context, following the procedure described in (Zhou et al., 2018) and also used in (Adams et al., 2023). Specifically, we greedily select source sentences until the ROUGE-1 F1 score between the set of selected source sentences and the summary stops increasing. The resulting set of source sentences forms a proxy of the visible salient input information being rephrased by the model when summarizing. We split each truncated source document into 10 bins of the same number of sentences, and map each aligned source sentence to its bin. Note that bins are not directly comparable across models, as context length varies across models.
As we can see in Fig. 2, sentences from the first 10% or last 10% of the input context are much more represented than others. A clear U-shape emerges on PubMed and SummScreenFD for all LLMs. This is intriguing knowing that Llama-2 LLMs and the other two LLMs have different context window lengths, and the last 10% of each context window may contain content of varying saliency. In other words, LLMs seem to be mostly re-phrasing information from the beginning or the end of their context window.
3.3 RQ3: Does LLMs performance depend on the position of salient information?
Results from the last experiment raise the question of whether LLMs’ summarization performance changes depending on where salient information is located within the input. As an approximation for salient information, we consider the alignment between summary sentences and source sentences like in Fig. 2, but this time using the reference summaries. Each reference summary is mapped to the source sentences it maximizes ROUGE-1 F1 against, which may be scattered across the whole source. We convert each source sentence to its cumulative word count from the beginning of the source, and take the average as an approximation of the mean position of salient information within the source. We keep data points with mean salient position fitting within the LLM context window.
We examine performance changes with regard to this salient position. To do so, we compute the Spearman correlation coefficient between salient position and each evaluation metric in Table 2. A high absolute Spearman value means that summary quality (as measured by this metric) can change (and deteriorate) with the position of important information within the context. There are several takeaway findings from this Table. First, we notice that on standard-length datasets, reference-based evaluation metrics are negatively correlated to position of salient information. The correlation is only moderate, yet remarkably consistent across datasets (except Reddit-TIFU) and models. This is surprising, since such datasets fit entirely in context and are not affected by truncation. In contrast, reference-free metrics show either no significant or positive correlation to information position. For long-input datasets, the negative trend for reference-based metrics is confirmed. On these lengthy datasets, SummaC and GPT-3.5 tend to switch from positive to negative correlation, especially for Xgen-7B and Mistral-7B. We highlight that since GPT-3.5 itself is affected by the middle-curse from Liu et al. (2023a), it may not accurately evaluate summarization when salient content lays in the middle of the context. In light of these results, we conclude that LLMs’ summarization performance is sensitive to the position of salient information in the context window.
4 Analysis
Figure 3: Multi-document summarization performance on Multi-XScience (top row) and Multi-News (bottom row) when a unique relevant document is used, and its position is varied (x-axis). Dashed horizontal lines correspond to the random baseline.
Figure 4: Fine-grained evaluation of multi-document summarization on Multi-News with GPT-3.5 when varying the position of a unique relevant input document.
4.1 How is information in the middle treated?
Previous experiments show that LLMs place more emphasis on the beginning and the end of their context. We now narrow down on how LLMs treat the middle. To remove the effect of spread of salient information, we perform two controlled experiments in multi-document summarization. This setup enables us to shuffle the order of the input, which is not realistic for the single-document setup as it would break coherence. We only consider data points with the same number $k$ of documents: $k=7$ documents for Multi-XScience ($n=329$), and $k=5$ documents for Multi-News ($n=219$).
In the first experiment, we vary the position of salient information throughout the input. We keep a single document (the abstract of the query paper on Multi-XScience, and the document with the highest BERTScore with the reference on Multi-News), and place it at position $j$ for $j\in\{1,\ldots,k\}$, using $k-1$ documents from a random data point for the other slots. The single relevant document is accompanied by a [RELEVANT] header, while the other documents have an [IRRELEVANT] header, and we prompt the LLM to only summarize the relevant document. For reference-free evaluation metrics, we use the single relevant document as source. In this experiment and the following, we also include a random baseline of shuffled inputs and model predictions. In Fig. 3, we see a noticeable drop in performance for all metrics when the salient document is not in the first or final position. Flan-UL2 seems to focus on the end of the context, Xgen-7B and Mistral-7B on the beginning, and Llama-2 models on both. Performance can fall quite below random range, especially for SummaC, confirming the worrying trend from Liu et al. (2023a).
A more fine-grained analysis with GPT-3.5 in Fig. 4 evaluating specific attributes (following the method in Adams et al. (2023)) reveals more details. Coherence and quality remain high and stable. In other words, the text outputs are of good quality. But for informativeness and attributability, the U-shape appears again: it shows that the LLMs are struggling to generate content sticking to the document inserted in the middle.
Dataset Model Input documents R-2 BS A3CU SummaC GPT-3.5
Multi-X Llama-2-7B All 7 4.64 82.83 5.88 54.56 4.17
First + last 4.62 82.82 5.78 47.00 4.50
First + 5 random + last 4.43 82.64 5.37 43.25 4.01
Random 1.22 81.14 3.08 30.30 1.14
[…]
First + last 4.73 82.86 5.76 43.74 4.53
First + 5 random + last 4.61 82.80 5.72 46.42 4.31
Random 1.16 81.16 3.22 28.38 1.42
[…]
First + last 5.01 82.73 5.86 49.08 4.45
First + 5 random + last 3.89 82.16 5.03 55.29 3.01
Random 1.25 80.74 3.21 27.77 1.29
[…]
First + last 5.12 82.67 6.15 60.91 4.76
First + 5 random + last 4.45 82.40 5.35 58.19 4.06
Random 1.25 80.64 3.13 25.97 1.18
Multi-N Llama-2-7B All 5 10.76 85.04 19.06 60.09 4.00
First + last 9.50 84.43 15.88 54.52 3.80
First + 3 random + last 7.57 83.36 12.39 50.35 2.94
Random 1.52 80.22 3.11 33.34 1.92
[…]
First + last 9.55 84.58 16.99 49.84 3.73
First + 3 random + last 8.27 83.79 14.84 50.09 3.18
Random 1.55 80.05 3.17 32.80 1.98
[…]
First + last 7.82 83.27 14.18 51.59 3.60
First + 3 random + last 6.30 81.85 11.66 49.02 2.66
Random 1.52 78.99 3.48 37.62 1.86
[…]
First + last 9.11 83.69 14.99 67.14 3.51
First + 3 random + last 6.59 81.66 12.50 52.14 2.45
Random 1.57 79.21 3.30 37.18 1.80
Table 3: Performance in multi-document summarization on Multi-XScience (7 documents) and Multi-News (5 documents) when infilling the middle of the context window with random documents. R-2 is ROUGE-2, BS refers to BERTScore.
Figure 5: Reference-based evaluation on the MiddleSum dataset. We also report (gray bars) performance achieved by uniformly sampling subsets of the same size as MiddleSum from the original datasets, alongside bootstrapping variance (black lines).
In the second experiment, we take the opposite approach, and put salient information at the beginning and the end, while the middle of the prompt is filled with noise. We keep the first and last documents, and fill the $k-2$ middle ones with random documents. We also run a baseline just using the first and last documents as input, expected to be close to the result with random documents in between. As displayed in Table 3, filling with random noise between the first and last document (which amounts to a prompt mostly irrelevant to the reference) leads to a moderate drop in performance. For instance, on Multi-XScience, with 5 random documents between the first and last, Llama-2-13B maintains 4.31 as GPT-3.5 score compared to 4.35 when using all 7 documents.
We conclude from these two experiments that LLMs can focus on the beginning and/or the end of their input, but largely ignore the middle. The U-shape or middle curse from Liu et al. (2023a) also applies to abstractive summarization.
4.2 Can we alleviate the middle curse?
To evaluate loss of performance due to the middle curse in a natural setup, we subsample data points from each of the 5 long-input summarization datasets. We obtain sentences from the (untruncated) source aligned with the reference summary, following the procedure from Section 3.2. Only data points where the start index of the earliest aligned source sentence is at least 1,200 words, are kept, ensuring no salient information at the start. We randomly sample 50 data points from each of Arxiv, PubMed, GovReport and Multi-News, and 25 from SummScreenFD, forming an evaluation dataset of 225 samples which we name MiddleSum. [^2] We will release it alongside our code
We evaluate LLMs on MiddleSum, keeping only reference-based evaluation as the dataset is built using saliency with regard to the reference. As expected, in Fig. 5 we see that LLMs perform noticeably worse on MiddleSum (green bars) as compared to the full set (gray bars), confirming that MiddleSum is a more challenging task.
We benchmark alternative inference methods on MiddleSum: hierarchical summarization and incremental summarization, both of which are explored in the concurrent work of Chang et al. (2023). Namely, let us divide an input ${\bm{x}}$ of length $n$ into $k$ consecutive blocks of size at most $m$ (yielding $k=\left\lceil{\frac{n}{m}}\right\rceil$): ${\bm{x}}=({\bm{x}}_{1},\ldots,{\bm{x}}_{k})$.
Hierarchical summarization consists in summarizing each block and then summarizing the concatenation of summaries:
${\bm{y}}_{i}$ $=\text{LLM}({\bm{x}}_{i})\quad\forall i\in\{1,\ldots,k\}$ (1)
${\bm{y}}$ $=\text{LLM}({\bm{y}}_{1},\ldots,{\bm{y}}_{k})$ (2)
Incremental summarization consists in updating a summary of the text so far with content from the current text block (we have ${\bm{y}}_{0}=\emptyset$):
$\forall i\in\{1,\ldots,k\},{\bm{y}}_{i}=\text{LLM}({\bm{y}}_{i-1},{\bm{x}}_{i})$ (3)
In both methods, the final output is ${\bm{y}}$. Noting $l$ the output length, standard inference has complexity in $\mathcal{O}(l.n^{2})$, while both alternative methods have complexity in $\mathcal{O}(k.l.m^{2})=\mathcal{O}(l.n.m)$, which is lower. For both methods and all models, we use a block size $m$ of 1,500 words (roughly 2,000 tokens), and preserve coherence by ending blocks at the earliest end of sentence reaching the length.
Results are shown in Fig. 5 (blue and purple bars), with statistics, bigrams analysis and results split over datasets in Appendix E. Both methods show promising results, notably on Mistral-7B for which they improve performance significantly. Across domains, hierarchical and incremental inference are very successful on scientific publications, which we hypothesize is due to the natural division in structured sections of such inputs. Yet, they seem to harm summaries on the other domains.
4.3 Is scaling context length really useful?
Figure 6: Long-input summarization performance on GovReport with 4 models and all 5 metrics. X-axis represents the truncated maximum source length. Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.
Experiments from Section 4.1 confirm that LLMs struggle to summarize information contained in the middle of their context window. This poses issues for long-input summarization: after the initial part with high saliency, important information becomes sparser, and at the same time LLMs processing capability weakens. To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6. We use our longest context LLMs Xgen-7B and Mistral-7B, as well as two LLMs extending Llama-2-7B context window with position interpolation (Chen et al., 2023), a method gaining traction as an efficient way to scale LLMs’ context window. We use two LLMs applying position interpolation: Vicuna-7B-1.5-16k [^3] In HuggingFace: lmsys/vicuna-7b-v1.5-16k, and Llama-2-7B-32k [^4] In HuggingFace: togethercomputer/LLaMA-2-7B-32K, with context of 16k tokens and 32k tokens, respectively.
Results for GovReport in Fig. 6 (and Appendix F for Arxiv and SummScreen) confirm our intuition: all metrics plateau or even decrease (see Mistral-7B) from 4k context window upwards. Two conflicting forces are at play when increasing length: giving more information to the model helps it retrieving key elements to make a summary, while reasoning over a longer context is more challenging. Yet, such a drop for Xgen-7B and Mistral-7B at 8k inference length is concerning. Both position interpolated models show more robustness ; thus these patterns may vary a lot across LLMs. Our results suggest that in the current LLMs inference and evaluation framework, there is no need to exceed 4k tokens in the context window.
5 Related Work
Summarization with LLMs
It is widely acknowledged that LLMs have propelled forward abstractive summarization research (Pu et al., 2023), with their summaries being highly rated by human annotators (Goyal et al., 2022; Zhang et al., 2023b). Liu et al. (2023b) proposes to train smaller models like BART (Lewis et al., 2020) or BRIO (Liu et al., 2022b) with contrastive learning using LLMs like ChatGPT as evaluator providing signal on which generated summary candidate is better. Summary chain-of-thought designs a custom chain-of-thought method which first prompts the LLM to list important facts, then integrates these facts into a coherent summary (Wang et al., 2023c). SummIt utilizes ChatGPT to iteratively write then refine summaries given feedback from an evaluator LLM (Zhang et al., 2023a). Chain-of-density gradually makes GPT-4 generated summaries contain more and more entities while keeping length budget constant, creating more informative albeit a bit less readable summaries (Adams et al., 2023). Ravaut et al. (2022) noticed that data points with higher compression are generally harder to summarize with pre-trained models.
Position bias in LLMs
Sun et al. (2021) showed that for Transformer-based models, most recent tokens play a greater role compared to older tokens for next-token prediction. It was later found that for in-context learning, the order of examples within the prompt impacts GPT-3’s performance Liu et al. (2021); Lu et al. (2021). Reliance on positional information affects LLMs capabilities in arithmetic (Shen et al., 2023b), in multiple-choice question-answering (Zheng et al., 2023; Pezeshkpour and Hruschka, 2023), and as text generation evaluators (Wang et al., 2023b) ; making it hard to rank LLMs (Alzahrani et al., 2024). Liu et al. (2023a) was the first to show that LLMs’ performance weakens in the middle of the prompt, yet, how LLMs make use of their full context window remains poorly understood. The passkey retrieval evaluation, which consists in prompting the LLM to recall a complex string or long number inserted in its prompt, is becoming popular recently as a way of verifying LLM’s processing capability at each position Liu et al. (2023a); Jiang et al. (2024). However, this task does not measure position bias on complex, abstract reasoning tasks like summarization. A line of work attempts to solve the middle curse through compressing the prompt Jiang et al. (2023b, c), with very promising results albeit at the cost of prompt fluency. Another approach marginalizes results over different permutations of the input to suppress dependency on input order Tang et al. (2023). Concurrent work to ours also finds that in zero-shot summarization, LLMs tend to prefer lead content Chhabra et al. (2024).
6 Conclusion
Behind the recent hype around LLMs and their amazing instruction following and content generation capacities, our study showcases a major weakness in abstractive summarization: LLMs suffer from the middle curse and struggle to use information in the middle of their context window. LLMs do not make a consistent use of their context window as they mostly look at the beginning and (to a lesser extent) the end, which at first glance may be hidden by the prevalent lead bias in summarization datasets. Extending context window beyond 4k tokens, which has been an intense area of focus lately, is not justified in the current inference and evaluation setup in abstractive summarization. We benchmarked two alternative inference methods (hierarchical inference and incremental inference) on MiddleSum, an evaluation subset designed to showcase the middle curse. Despite promising results, especially on scientific paper datasets, these methods are far from a silver bullet to the middle curse. We call for a better evaluation of LLMs, which accounts for the salient spans of the source which are effectively being processed.
Limitations
Our work only considers open-source LLMs for summary generation and ignores closed-source LLMs such as the popular OpenAI’s GPT-3.5 and GPT-4, or Anthropic’s Claude. We made this decision to advocate for openness in LLM research ; yet we acknowledge that it would be interesting to also investigate properties of summaries generated by these paying API LLMs.
Another limitation lays in the saliency estimation. We approximate salient content in the source through maximizing ROUGE-1 overlap with summary sentences. Other methods are also well-suited for this task, albeit at greater computational cost ; for instance semantic similarity through BERTScore or BARTScore ; or saliency estimation through a LLM in zero-shot.
Lastly, we can only evaluate a finite number of LLMs, and we settled for the evaluation of 5 recent and popular open-source LLMs. Findings may change as LLMs undergo changes and improvements in their training process.
Appendix A Statistics
Dataset Domain # Docs # Data points # Sentences # Words # Tokens
Train Val Test Doc. Summ. Inf. Doc. Summ. Doc. Summ. Max gen.
CNN/DM (Hermann et al., 2015) News 1.00 287,113 13,334 11,490 33.37 3.79 3 773.23 57.75 994.56 84.47 192
XSum (Narayan et al., 2018) News 1.00 204,045 11,332 11,334 19.11 1.00 1 433.05 23.19 566.79 31.63 64
Reddit-TIFU (Long) (Kim et al., 2019) Social Media 1.00 33,701 4,214 4,221 22.21 1.45 2 444.20 23.37 532.18 29.82 128
SAMSum Kim et al. (2019) Dialogue 1.00 14,732 818 819 8.96 2.01 2 126.93 23.12 175.54 29.69 128
Multi-XScience (Lu et al., 2020) Science 5.06 30,369 5,066 5,093 30.55 4.86 5 773.36 120.65 965.99 157.77 384
[…]
PubMed (Cohan et al., 2018) Science (medical) 1.00 119,924 6,633 6,658 101.61 7.59 7 3,142.92 208.03 4,602.62 324.97 512
GovReport (Huang et al., 2021) Legal 1.00 17,517 973 973 282.86 23.14 22 8,363.22 649.01 11,025.02 879.10 768
SummScreenFD (Chen et al., 2022) TV Transcripts 1.00 3,673 338 337 727.06 5.26 5 7,618.20 123.34 10,067.39 157.44 512
Multi-News (Fabbri et al., 2019) News 2.73 44,972 5,622 5,622 79.02 9.88 10 2,101.49 256.55 2,998.52 324.29 512
Table 4: Statistics on the 10 datasets used for experiments. Doc. is the source document, Summ. the ground-truth summary, Inf. refers to the number of desired sentences to be in the summary prompted to each LLM during inference. Statistics are computed on the entire test set. # Tokens is calculated with Llama-2’s tokenizer. Max gen. is the maximum tokens size that we set when decoding summaries. Underlined test sizes correspond to datasets where we subsample randomly 1,000 test data points for evaluation.
In Table 4, we include statistics on each of the abstractive summarization datasets under consideration. We use the non-anonymized version for CNN/DM (See et al., 2017). For Reddit-TIFU, we use the Long subset, and SummScreenFD, is the ForeverDreaming (FD) subset of SummScreen. GovReport and SummScreenFD are part of the long-input benchmarks Scrolls (Shaham et al., 2022) and ZeroScrolls (Shaham et al., 2023).
In Fig. 7, we illustrate how much of the source document(s) is visible with a 4k context window (Llama-2).
Figure 7: Fraction of the source which fits into the context window, for several context window lengths with Llama-2 tokenization. The black dashed lines correspond to Llama-2 context window length of 4k tokens. On standard-length datasets, 4k is enough to access 100% of all source documents ; but on the long-input datasets such as GovReport or SummScreenFD, such a context window may not even fit 50% of the source.
Appendix B GPT-3.5 Evaluation
Model Metric
CNN/DM
XSum
Reddit-TIFU
SAMSum
Multi-X
AVG
Arxiv
PubMed
GovReport
SummScreenFD
Multi-N
AVG
SOTA ROUGE-2 24.17 27.09 11.13 29.88 4.60 19.37 21.93 23.26 30.90 10.70 13.60 20.07
[…]
ROUGE-2 20.28 22.74 8.61 28.21 3.04 16.58 9.37 7.42 4.76 4.33 7.79 6.73
BERTScore 88.05 91.94 87.42 92.29 81.87 88.31 83.82 83.07 83.53 84.85 84.97 84.05
A3CU 32.69 32.11 16.89 49.48 5.98 27.43 14.79 13.83 12.00 8.37 16.99 13.20
SummaC 69.96 24.27 35.76 30.19 57.98 43.63 67.56 60.96 73.80 56.00 76.80 67.02
GPT-3.5 3.16 3.52 1.61 2.92 3.23 2.89 3.11 3.16 2.46 2.11 3.36 2.84
[…]
ROUGE-2 14.16 7.27 4.17 15.53 4.87 9.20 13.84 12.89 16.22 5.36 12.37 12.14
BERTScore 87.32 87.47 85.87 89.95 83.32 86.79 83.84 82.82 85.28 85.41 85.63 84.60
A3CU 29.04 14.18 12.15 35.64 6.39 19.48 16.78 16.60 17.23 9.66 22.23 16.50
SummaC 35.58 25.24 26.38 24.56 49.39 32.23 53.22 51.82 70.47 39.01 57.49 54.40
GPT-3.5 4.10 4.24 2.83 3.61 4.42 3.84 4.21 4.19 3.43 2.71 3.91 3.69
[…]
ROUGE-2 14.10 8.61 4.22 14.16 5.29 9.28 13.52 15.24 17.28 5.62 12.58 12.85
BERTScore 87.40 87.94 85.85 89.45 83.58 86.84 83.88 84.24 85.29 85.42 85.84 84.93
A3CU 29.57 16.30 12.94 34.19 7.29 20.06 16.44 19.21 17.01 10.30 23.09 17.21
SummaC 33.83 24.07 25.76 24.81 41.70 30.03 55.09 56.00 76.44 38.74 53.12 55.88
GPT-3.5 4.12 4.34 2.91 3.45 4.45 3.85 4.06 4.13 3.66 2.69 3.89 3.69
[…]
ROUGE-2 14.55 6.00 3.98 14.51 5.54 8.92 12.31 13.99 14.68 4.24 11.07 11.26
BERTScore 87.07 87.12 85.84 89.53 83.42 86.60 83.07 82.87 83.94 83.91 84.95 83.75
A3CU 27.88 12.75 12.56 33.18 7.45 18.76 15.28 18.79 15.65 8.44 21.33 18.90
SummaC 52.25 37.95 28.40 25.63 44.36 37.72 53.28 60.56 65.22 42.03 56.45 55.51
GPT-3.5 3.82 3.97 2.78 3.52 4.37 3.69 3.96 3.99 2.78 2.42 3.58 3.35
[…]
ROUGE-2 16.37 7.05 4.34 14.66 5.57 9.60 9.77 14.32 11.36 3.11 12.61 10.23
BERTScore 87.50 87.45 85.71 89.78 83.16 86.72 81.44 82.85 82.43 81.46 85.07 82.65
A3CU 30.60 13.14 13.08 32.21 6.96 19.20 12.66 16.31 14.92 8.40 22.22 14.90
SummaC 53.67 24.79 30.51 26.82 62.49 39.66 57.81 69.03 67.07 35.76 68.50 59.63
GPT-3.5 3.92 4.30 2.73 3.63 4.47 3.81 2.83 3.63 2.01 1.88 3.60 2.79
Table 5: Performance achieved by the 5 LLMs on each dataset for 5 automatic metrics. # sents is the average number of sentences in generated summaries. Multi-X is short for Multi-XScience, Multi-N is Multi-News dataset, AVG columns represent the average over standard-length and long-input datasets, respectively. SOTA numbers are taken from(Xie et al., 2023) on CNN/DM, from (Zhao et al., 2022b) XSum, Reddit-TIFU and SAMSum, from (Pang et al., 2022) for Arxiv and PubMed, from (Xiong et al., 2022) for GovReport and SummScreenFD, from (Xiao et al., 2022) for Multi-News and Multi-XScience. Due to a lack of reported results for other metrics, we only include ROUGE-2 scores for SOTA models. Best scores (outside of SOTA) are in bold.
To evaluate LLM-generated summaries with GPT-3.5, we use the following prompt template:
Score the following summary generated by another system given the source on a scale from 1 to 5 with regards to overall general summary quality. 1-point indicates a low quality summary, and 5 points a very high quality summary. A high quality summary is grammatical, fluent, informative, relevant, coherent and factually consistent with the source. Let’s think step-by-step and just output the score.
Source:
[source]
Instruction:
Summarize the above text in [n] sentences.
Summary:
[summary generated by the LLM to evaluate]
Your score:
When evaluating for the specific aspect of informativeness, the first paragraph becomes:
Score the following summary generated by another system given the source on a scale from 1 to 5 with regards to how informative the summary is. 1 point indicates a not informative summary, and 5 points a very informative summary. An informative summary captures the important information in the article and presents it accurately and concisely. Let’s think step-by-step and just output the score.
When evaluating for the specific aspect of overall quality, the first paragraph becomes:
Score the following summary generated by another system given the source on a scale from 1 to 5 with regards to its quality. 1 point indicates a low quality summary, and 5 points a very high quality summary. A high quality summary is comprehensible and understandable. Let’s think step-by-step and just output the score.
Model CNN/DM XSum Reddit-TIFU SAMSum Multi-XScience Arxiv PubMed GovReport SummScreenFD Multi-News
Flan-UL2 0.000 0.012 0.000 0.017 $0$
Llama-2-7B 0.000 0.000 0.000 0.490 0.000 0.000 0.000 0.000 0.000 0.000
Llama-2-13B 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Xgen-7B 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Mistral-7B 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
Table 6: P-value of a 2-sample Kolmogorov-Smirnov test between the position distribution of bigrams in LLM-generated summaries and bigrams in reference summaries. We round numbers to 3 decimals. Numbers in gray correspond to non-significant differences (p-value above 0.05).
Dataset Domain # Docs # Data points # Sentences # Words # Tokens
Doc. Summ. Doc. Summ. Doc. Summ.
Arxiv Science 1.00 50 299.36 5.84 7,605.52 165.80 10,846.00 226.22
PubMed Science (medical) 1.00 50 157.60 6.96 5,090.84 203.04 7,783.02 329.12
GovReport Legal 1.00 50 445.64 22.80 13,308.70 656.10 17,462.90 883.10
SummScreen TV transcripts 1.00 25 762.24 3.56 9,732.44 88.60 12,878.08 115.00
MultiNews News 3.18 50 211.34 9.86 5,939.32 269.84 8,130.78 336.16
[…]
Table 7: Statistics on the MiddleSum evaluation dataset, breaking down on each domain. Doc. is the source document, Summ. the ground-truth summary. # Tokens is calculated with Llama-2’s tokenizer.
When evaluating for the specific aspect of coherence, the first paragraph becomes:
Score the following summary generated by another system given the source on a scale from 1 to 5 with regards to its coherence. 1 point indicates an incoherent summary, and 5 points a very coherent summary. A coherent summary is well-structured and well-organized. Let’s think step-by-step and just output the score.
When evaluating for the specific aspect of attributability, the first paragraph becomes:
Score the following summary generated by another system given the source on a scale from 1 to 5 with regards to how attributable it is. 1 point indicates a not very attributable summary, with many hallucinations, and 5 points a summary very attributable to the source, consistent with the source. In a very attributable summary, all the information is fully attributable to the source. Let’s think step-by-step and just output the score.
Appendix C Baseline Performance
(a) Llama-2-7B
(b) Llama-2-13B
(c) Xgen-7B
(d) Mistral-7B
Figure 8: Relative location of summary bigrams within the source on MiddleSum for 4 LLMs.
In Table 5, we report zero-shot performance with the 5 LLMs described in Section 2. We note that for standard-length datasets, Flan-UL2 is dominating, while for long-context ones, Llama-2-13B takes the lead.
Appendix D Statistical Significance on RQ1
In Table 6, we run a 2-sample Kolmogorov-Smirnov statistical significance test to compare the bigrams position distribution of LLMs with the reference summaries.
Appendix E More Results on MiddleSum
In Table 7, we show statistics on the MiddleSum evaluation dataset.
In Figs. 7(a), 7(b), 7(c) and 7(d), we repeat the salient bigrams analysis from Fig. 1 for each of the four long-input LLMs on MiddleSum. We note that both hierarchical and incremental inference notably decrease reliance on lead bigrams compared to standard inference.
In Figs. 9, 10, 11, 12 and 13, and in Table 8 (exact numbers), we show reference-based evaluation on each of the 5 subsets of MiddleSum.
Figure 9: Reference-based evaluation on the Arxiv subset of MiddleSum dataset.
Figure 10: Reference-based evaluation on the PubMed subset of MiddleSum dataset.
Figure 11: Reference-based evaluation on the GovReport subset of MiddleSum dataset.
Figure 12: Reference-based evaluation on the SummScreenFD subset of MiddleSum dataset.
Figure 13: Reference-based evaluation on the Multi-News subset of MiddleSum dataset.
Appendix F More Results on Long-Input Tasks
In Figs. 14 and 15, we conduct the analysis on increasingly longer input context from Fig. 6 on Arxiv and SummScreen, respectively.
Figure 14: Long-input summarization performance on Arxiv with 4 models and all 5 metrics. X-axis represents the truncated maximum source length. Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.
Figure 15: Long-input summarization performance on SummScreen with 4 models and all 5 metrics. X-axis represents the truncated maximum source length. Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.
Model Metric Inference MiddleSum (MS) MS/Arxiv MS/PubMed MS/GovReport MS/SummScreen MS/Multi-News
Llama-2-7B ROUGE-2 Standard 10.96 12.62 10.97 13.26 4.07 10.43
Hierarchical 11.33 14.63 13.36 13.51 4.85 7.06
Incremental 12.56 14.43 13.54 17.26 5.19 8.68
[…]
Hierarchical 84.13 83.86 83.74 84.26 85.23 84.10
Incremental 84.26 83.73 83.80 84.67 85.70 84.14
[…]
Hierarchical 13.21 15.64 15.14 12.26 9.71 11.57
Incremental 12.88 15.34 14.45 10.51 10.24 12.55
Llama-2-13B ROUGE-2 Standard 11.07 11.68 11.63 13.56 5.09 10.38
Hierarchical 10.24 13.45 13.26 10.21 4.93 6.71
Incremental 11.90 12.53 13.84 17.34 5.06 7.33
[…]
Hierarchical 83.17 83.74 83.75 80.07 85.03 84.20
Incremental 83.43 82.45 83.45 84.96 83.24 82.95
[…]
Hierarchical 12.85 16.28 14.11 10.40 8.36 12.86
Incremental 12.47 14.64 14.80 11.75 10.76 9.54
Xgen-7B ROUGE-2 Standard 8.92 11.94 9.34 10.02 4.53 6.57
Hierarchical 9.87 13.06 11.37 11.77 2.79 6.83
Incremental 9.11 10.97 10.55 13.16 3.71 4.44
[…]
Hierarchical 82.86 82.73 82.56 82.67 83.79 83.03
Incremental 82.49 81.94 81.82 83.11 83.39 80.83
[…]
Hierarchical 11.74 14.55 12.61 10.76 7.33 11.25
Incremental 11.51 12.73 13.15 9.13 8.97 9.69
Mistral-7B ROUGE-2 Standard 7.16 9.28 9.56 5.86 2.19 6.41
Hierarchical 10.78 11.67 13.71 13.36 4.76 7.39
Incremental 10.15 11.36 12.66 12.89 3.46 7.02
[…]
Hierarchical 83.10 82.56 83.00 83.81 83.90 82.62
Incremental 82.25 81.92 82.39 83.67 83.14 80.56
[…]
Hierarchical 11.15 12.15 14.21 10.62 7.48 9.47
Incremental 10.43 14.25 11.00 8.90 7.44 9.05
Table 8: Reference-based results for all models and inference methods on MiddleSum, breaking down by subset. The best number across inference methods is in bold.
Appendix G Software
We use the following Python libraries, all open source:
  • numpy, version 1.24.3
  • torch, version 2.0.1
  • scikit-learn, version 1.0.2
  • sentencepiece, version 0.1.97
  • nltk, version 3.8.1
  • spacy, version 3.6.0
  • scipy, version 1.10.1
  • rouge-score, version 0.1.2
  • bert-score, version 0.3.13
  • summac, version 0.0.03
  • tiktoken, version 0.4.0
  • openai, version 0.28.0
  • huggingface-hub, version 0.17.2
  • datasets, version 2.14.5
  • accelerate, version 0.21.0
  • tokenizers, version 0.14.1
  • transformers, version 4.34.0

### v4

@@
 In this paper, we conduct the first comprehensive study on context utilization and position bias in summarization.
-Our analysis encompasses 5 LLMs, 10 datasets, and 5 evaluation metrics.
-We introduce a new evaluation benchmark called MiddleSum on the which we benchmark two alternative inference methods to alleviate position bias: hierarchical summarization and incremental summarization [^1] We will share all code and data upon publication..
+Our analysis encompasses 6 LLMs, 10 datasets, and 5 evaluation metrics.
+We introduce a new evaluation benchmark called MiddleSum on the which we benchmark two alternative inference methods to alleviate position bias: hierarchical summarization and incremental summarization [^1] Our code and data can be found here: https://github.com/ntunlp/MiddleSum..
 1 Introduction
@@
 (2022); Chung et al.
-(2022), LLMs made a major leap forward in conditional (prompted) content generation, and can generate satisfying outputs without the need to finetune on a task specifically.
-In abstractive summarization specifically, this approach has arguably opened a new paradigm; summaries generated by LLMs are highly fluent, grammatical and relevant (Goyal et al., 2022).
+(2022), LLMs made a major leap forward in conditional (prompted) content generation, and can generate satisfying outputs without the need to fine-tune on a specific task.
+In abstractive summarization specifically, this approach has arguably opened a new paradigm: summaries generated by LLMs are highly fluent, grammatical and relevant (Goyal et al., 2022).
 Despite noticeably lower scores on automatic metrics such as ROUGE (Lin, 2004) or BERTScore (Zhang et al., 2019), summaries generated by LLMs are largely preferred by humans over summaries from state-of-the-art fine-tuned models like BRIO (Liu et al., 2022b, 2023c).
-In fact, on XSum, GPT-3.5 summaries are even on par with re-annotated human-written summaries, and much better than the dataset ground-truth, according to human evaluators (Zhang et al., 2023b).
-LLMs also show promising capability in evaluating summaries generated by other systems, including other LLMs (Fu et al., 2023; Luo et al., 2023; Shen et al., 2023a).
+In fact, on XSum, GPT-3.5 summaries are even on par with re-annotated human-written summaries, and much better than the dataset’s original ground-truth, according to human evaluators (Zhang et al., 2023b).
+LLMs also show promising capability in evaluating summaries generated by other systems, including LLMs (Fu et al., 2023; Luo et al., 2023; Shen et al., 2023a).
 Despite this success, a few major technological bottlenecks remain with LLMs, including the maximum length of their context window.
@@
 Several techniques were proposed to extend the context window, including ALiBi (Press et al., 2021), LeX (Sun et al., 2022), position interpolation (Chen et al., 2023) and YaRN (Peng et al., 2023).
-While some of them claim up to 128k tokens processing capacity (Peng et al., 2023), it remains unclear how much such methods help on long-context summarization.
+While some of them claim up to 100k+ tokens processing capacity (Peng et al., 2023), it remains unclear how much such methods help on long-context summarization.
 Scaling up context length would only succeed if a key question gets addressed first: do LLMs make proper use of their entire context?
@@
 Unlike in question-answering, mapping facts in the output to a specific snippet in the source is not straightforward in abstractive summarization, due to the high-level of re-phrasing and compression.
-We conduct a large-scale study with 5 open-source LLMs, 10 datasets covering many aspects of summarization, and 5 highly diverse automatic metrics.
+We conduct a large-scale study with 6 LLMs, 10 datasets covering many aspects of summarization, and 5 highly diverse automatic metrics.
 Our contributions are threefold:
@@
 Models
-We experiment with several popular high-performing open-source instruction-tuned LLMs.
-Instruction-tuning datasets such as Flan (Wei et al., 2021) include some of the datasets we study: CNN/DM, XSum, SAMSum and Multi-News.
-We analyze the following models through HuggingFace transformers (Wolf et al., 2020):
+We experiment with 6 popular and high-performing LLMs:
   • Flan-UL2 is a 20B parameters encoder-decoder model pre-trained on 1T tokens.
 It is based on the UL2 20B model (Tay et al., 2022), with the addition of Flan-T5 (Chung et al., 2022) instruction fine-tuning.
-The context window is 2k.
-  • Llama-2 (Touvron et al., 2023b) is a recently introduced powerful decoder-only model pre-trained on 2T tokens, ranging from 7B to 70B parameters, and with a 4k context window.
+The context window is 2k tokens.
+  • Llama-2 (Touvron et al., 2023b) is a recently introduced powerful decoder-only model pre-trained on 2T tokens, ranging from 7B to 70B parameters, and with a 4k tokens context window.
 We use the 7B and 13B models.
-  • Xgen-7B (Nijkamp et al., 2023) is a 7B decoder-only model pre-trained on up to 1.5T tokens. It supports 8k tokens context window.
+  • Xgen-7B (Nijkamp et al., 2023) is a 7B decoder-only model pre-trained on up to 1.5T tokens. It supports an 8k tokens context window.
   • Mistral-7B (Jiang et al., 2023a) is also an 8k-context 7B decoder-only model, with performance slightly better than Llama-2-13B.
+  • GPT-3.5 (gpt-3.5-turbo-0125), which is known to be good at generating summaries and has a 16k tokens context window.
+[^2] To reduce cost, we subsample 300 data points when using GPT-3.5.
+Our total API cost is inferior to 300 USD.
 Input length # Documents In Flan?
@@
 Table 1: Summarization datasets under study. In standard length datasets, the context and summary fit within a 2k tokens LLM context window.
+We analyze the open-source models through HuggingFace transformers library (Wolf et al., 2020), and use the OpenAI API for GPT-3.5.
 We use the instruction-tuned (or chat) checkpoints for Llama-2, Xgen-7B and Mistral-7B.
+Note that popular instruction-tuning datasets such as Flan (Wei et al., 2021) include some of the datasets we study: CNN/DM, XSum, SAMSum and Multi-News.
 To run inference, we use the following prompt: Read the following text and summarize it: [text].
@@
 Summary: where n is set to an average number of target sentences per dataset (see Appendix A).
-We infer all models in bfloat16 and sample summaries with top-k sampling (Fan et al., 2018) using $k=50$.
+We infer all models in bfloat16 and sample summaries with top-k sampling (Fan et al., 2018) using $k=50$ and temperature $T=0.3$.
 Evaluation Measures
-Summarization evaluation is especially challenging in the LLM era, as most automatic metrics poorly correlate with human preferences (Goyal et al., 2022; Liu et al., 2022a).
+Summarization evaluation is especially challenging in the LLM era, as most automatic metrics poorly correlate with human preferences (Goyal et al., 2022; Liu et al., 2023b).
 To get a broad picture of performance, we evaluate with metrics as diverse as possible.
-First, we consider reference-based metrics: ROUGE-2 (Lin, 2004), which measures bigram overlap, BERTScore (Zhang et al., 2019), which measures semantic similarity with BERT (Devlin et al., 2018) embeddings, and A3CU (Liu et al., 2023c), which extracts facts in the form of Atomic Content Units (ACUs) Liu et al.
-(2022a), and checks the presence of ACUs between prediction and reference.
+First, we consider reference-based metrics: ROUGE-2 (Lin, 2004), which measures bigram overlap, BERTScore (Zhang et al., 2019), which measures semantic similarity with BERT (Devlin et al., 2019) embeddings, and A3CU (Liu et al., 2023c), which extracts facts in the form of Atomic Content Units (ACUs) Liu et al.
+(2023b), and checks the presence of ACUs between prediction and reference.
 As reference-free metrics, we include SummaC (Laban et al., 2022), a leading factual consistency evaluation metric relying on entailment scores between pairs of source and summary sentences.
-We also leverage GPT-3.5 (gpt-3.5-turbo-0125) as a summarization evaluator, which is proven to be a strong natural language generation evaluator (Wang et al., 2023a; Shen et al., 2023a; Jain et al., 2023).
+We also leverage GPT-3.5 again (still gpt-3.5-turbo-0125), this time as a summarization evaluator, which is proven to be a strong natural language generation evaluator (Wang et al., 2023a; Shen et al., 2023a; Jain et al., 2023).
 We prompt the model with the source and generated summary (which fits in GPT-3.5’s 16k context window) and ask to output a score on a likert scale from 1 to 5.
 We refer to Appendix B for the full prompt template.
-For GPT-3.5 evaluation, we subsample 300 data points per dataset in order to reduce costs.
+[^3] We also subsample 300 data points when using GPT-3.5 as an evaluator, to reduce API cost.
 We report the performance of LLMs on all 10 datasets, alongside a comparison to SOTA, in Appendix C.
-FLan-UL2 dominates on standard-length datasets, but Llama-2-13B has the upper hand on the long-input ones.
+FLan-UL2 dominates on standard-length datasets, but GPT-3.5 has the upper hand on the long-input ones.
 Performance itself is not our focus in this paper, but rather which position-related factors influence it.
-We discard Flan-UL2 on long-input datasets due to poor performance.
+We discard Flan-UL2 on long-input datasets due to very poor performance.
 Figure 2: Distribution of relative location of input context sentences aligned with sentences from summaries.
@@
 3.1 RQ1: Where in the source do LLMs take their information from?
-We investigate summaries generated by LLMs, and map them to specific parts of the input.
+We first investigate summaries generated by LLMs, and map them to specific parts of the input.
 Unlike in question-answering or extractive summarization, mapping salient information from a summary to the source is not trivial in abstractive summarization.
@@
 On XSum, Flan-UL2 closely matches the reference distribution, which we attribute to its better instruction tuning.
-Results in Appendix D confirm that bigram distribution for LLMs and references are statistically different (p-value of Kolmogorov-Smirnov test (Massey Jr, 1951) inferior to 0.001) in all cases except Flan-UL2 on XSum and SAMSum and Llama-2-7B on XSum.
+Results in Appendix D confirm that bigram distribution for LLMs and references are statistically different (p-value of Kolmogorov-Smirnov test (Massey Jr, 1951) inferior to 0.001) in all but 4 out of the 55 (dataset, LLM) setups: Flan-UL2 on XSum and SAMSum, Llama-2-7B on SAMSum and GPT-3.5 on SAMSum.
 We conclude that LLMs focus on contents at the beginning of the source document(s).
@@
 Mistral-7B -0.289 -0.031 0.006 -0.024 -0.052 -0.078 -0.270 -0.279 -0.585 -0.132 -0.324 -0.318
+GPT-3.5 -0.323 -0.027 -0.031 -0.097 0.088 -0.078 0.026 -0.093 -0.123 -0.061 -0.233 -0.097
 […]
@@
 Mistral-7B -0.278 -0.052 0.014 -0.108 -0.416 -0.168 -0.348 -0.367 -0.567 -0.356 -0.403 -0.408
+GPT-3.5 -0.280 -0.021 -0.025 -0.174 -0.247 -0.149 -0.087 -0.210 -0.175 -0.142 -0.262 -0.175
 […]
@@
 Mistral-7B -0.291 -0.110 0.004 -0.105 -0.119 -0.126 -0.160 -0.283 -0.305 -0.010 -0.283 -0.208
+GPT-3.5 -0.256 -0.022 -0.039 -0.216 0.141 -0.078 -0.064 -0.293 -0.201 -0.059 -0.264 -0.176
 SummaC Flan-UL2 -0.012 0.548 0.270 0.186 -0.035 0.191 _ _ _ _ _ _
@@
 Mistral-7B -0.045 0.515 0.149 0.069 0.154 0.128 -0.250 -0.103 -0.387 0.124 -0.010 -0.125
+GPT-3.5 0.156 0.590 0.444 0.180 0.008 0.276 0.058 0.237 0.061 -0.055 0.089 0.078
 […]
@@
 Mistral-7B -0.115 0.124 -0.133 0.036 -0.204 -0.058 -0.446 -0.322 -0.580 -0.188 -0.163 -0.342
+GPT-3.5 0.024 0.014 -0.108 0.131 -0.054 0.001 0.156 -0.008 -0.172 0.068 0.024 0.014
 Table 2: Spearman correlation coefficient between each LLM’s metric, and the mean position of salient information within the context window.
@@
 We align sentences in generated summaries to sentences in the context, following the procedure described in (Zhou et al., 2018) and also used in (Adams et al., 2023).
-Specifically, we greedily select source sentences until the ROUGE-1 F1 score between the set of selected source sentences and the summary stops increasing.
+Specifically, we greedily select source sentences (among the ones fitting in context window) until the ROUGE-1 F1 score between the set of selected source sentences and the summary stops increasing.
 The resulting set of source sentences forms a proxy of the visible salient input information being rephrased by the model when summarizing.
@@
 A clear U-shape emerges on PubMed and SummScreenFD for all LLMs.
-This is intriguing knowing that Llama-2 LLMs and the other two LLMs have different context window lengths, and the last 10% of each context window may contain content of varying saliency.
+This is intriguing knowing that the LLMs have different context window lengths, and the last 10% of each context window may contain content of varying saliency.
 In other words, LLMs seem to be mostly re-phrasing information from the beginning or the end of their context window.
+Results in Appendix E confirm an even stronger position bias with base models.
 3.3 RQ3: Does LLMs performance depend on the position of salient information?
@@
 (2023a), it may not accurately evaluate summarization when salient content lays in the middle of the context.
-In light of these results, we conclude that LLMs’ summarization performance is sensitive to the position of salient information in the context window.
+In light of these results, we conservatively conclude that LLMs’ summarization performance is sensitive to the position of salient information in the context window.
 4 Analysis
@@
 For reference-free evaluation metrics, we use the single relevant document as source.
-In this experiment and the following, we also include a random baseline of shuffled inputs and model predictions.
+We also include a random baseline of shuffled inputs and model predictions.
 In Fig.
 3, we see a noticeable drop in performance for all metrics when the salient document is not in the first or final position.
-Flan-UL2 seems to focus on the end of the context, Xgen-7B and Mistral-7B on the beginning, and Llama-2 models on both.
-Performance can fall quite below random range, especially for SummaC, confirming the worrying trend from Liu et al.
+Flan-UL2 seems to focus on the end of the context, Xgen-7B and Mistral-7B on the beginning, and Llama-2 models and GPT-3.5 on both.
+Performance can fall quite below random range, especially for reference-free metrics, confirming the worrying trend from Liu et al.
 (2023a).
@@
 4 evaluating specific attributes (following the method in Adams et al.
-(2023)) reveals more details.
+(2023), see Appendix B) reveals more details.
 Coherence and quality remain high and stable.
-In other words, the text outputs are of good quality.
-But for informativeness and attributability, the U-shape appears again: it shows that the LLMs are struggling to generate content sticking to the document inserted in the middle.
-Dataset Model Input documents R-2 BS A3CU SummaC GPT-3.5
-Multi-X Llama-2-7B All 7 4.64 82.83 5.88 54.56 4.17
-First + last 4.62 82.82 5.78 47.00 4.50
-First + 5 random + last 4.43 82.64 5.37 43.25 4.01
-Random 1.22 81.14 3.08 30.30 1.14
+In other words, the text outputs from the LLMs are always of good quality.
+But for informativeness and attributability, the U-shape appears again: it shows that the LLMs (even the powerful GPT-3.5) are struggling to generate content specifically sticking to the document inserted in the middle.
+Dataset Model Input documents R-2 BS A3CU SummaC GPT-3.5 %
+Multi-X Llama-2-7B All 7 4.64 82.83 5.88 54.56 4.17 100.00
+First + last 4.62 82.82 5.78 47.00 4.50 98.36
+First + 5 random + last 4.43 82.64 5.37 43.25 4.01 92.40
 […]
-First + last 4.73 82.86 5.76 43.74 4.53
-First + 5 random + last 4.61 82.80 5.72 46.42 4.31
-Random 1.16 81.16 3.22 28.38 1.42
+First + last 4.73 82.86 5.76 43.74 4.53 98.46
+First + 5 random + last 4.61 82.80 5.72 46.42 4.31 98.07
 […]
-First + last 5.01 82.73 5.86 49.08 4.45
-First + 5 random + last 3.89 82.16 5.03 55.29 3.01
-Random 1.25 80.74 3.21 27.77 1.29
+First + last 5.01 82.73 5.86 49.08 4.45 99.83
+First + 5 random + last 3.89 82.16 5.03 55.29 3.01 88.93
 […]
-First + last 5.12 82.67 6.15 60.91 4.76
-First + 5 random + last 4.45 82.40 5.35 58.19 4.06
-Random 1.25 80.64 3.13 25.97 1.18
-Multi-N Llama-2-7B All 5 10.76 85.04 19.06 60.09 4.00
-First + last 9.50 84.43 15.88 54.52 3.80
-First + 3 random + last 7.57 83.36 12.39 50.35 2.94
-Random 1.52 80.22 3.11 33.34 1.92
+First + last 5.12 82.67 6.15 60.91 4.76 99.80
+First + 5 random + last 4.45 82.40 5.35 58.19 4.06 90.59
 […]
-First + last 9.55 84.58 16.99 49.84 3.73
-First + 3 random + last 8.27 83.79 14.84 50.09 3.18
-Random 1.55 80.05 3.17 32.80 1.98
+First + last 4.75 83.04 6.05 39.81 4.69 96.42
+First + 5 random + last 4.37 82.90 5.67 43.26 4.63 95.63
+Multi-N Llama-2-7B All 5 10.76 85.04 19.06 60.09 4.00 100.00
+First + last 9.50 84.43 15.88 54.52 3.80 91.32
+First + 3 random + last 7.57 83.36 12.39 50.35 2.94 78.13
 […]
-First + last 7.82 83.27 14.18 51.59 3.60
-First + 3 random + last 6.30 81.85 11.66 49.02 2.66
-Random 1.52 78.99 3.48 37.62 1.86
+First + last 9.55 84.58 16.99 49.84 3.73 93.93
+First + 3 random + last 8.27 83.79 14.84 50.09 3.18 86.14
 […]
-First + last 9.11 83.69 14.99 67.14 3.51
-First + 3 random + last 6.59 81.66 12.50 52.14 2.45
-Random 1.57 79.21 3.30 37.18 1.80
+First + last 7.82 83.27 14.18 51.59 3.60 92.68
+First + 3 random + last 6.30 81.85 11.66 49.02 2.66 79.51
+[…]
+First + last 9.11 83.69 14.99 67.14 3.51 100.37
+First + 3 random + last 6.59 81.66 12.50 52.14 2.45 80.17
+[…]
+First + last 8.94 84.52 15.85 45.94 4.01 92.76
+First + 3 random + last 8.53 84.31 15.33 44.67 3.80 89.81
 Table 3: Performance in multi-document summarization on Multi-XScience (7 documents) and Multi-News (5 documents) when infilling the middle of the context window with random documents.
 R-2 is ROUGE-2, BS refers to BERTScore.
+% is the mean relative performance compared to the baseline with all documents.
 Figure 5: Reference-based evaluation on the MiddleSum dataset.
@@
 As displayed in Table 3, filling with random noise between the first and last document (which amounts to a prompt mostly irrelevant to the reference) leads to a moderate drop in performance.
-For instance, on Multi-XScience, with 5 random documents between the first and last, Llama-2-13B maintains 4.31 as GPT-3.5 score compared to 4.35 when using all 7 documents.
+For instance, on Multi-XScience, with 5 random documents between the first and last, Llama-2-13B maintains 98% of its performance, and reaches a GPT-3.5 score of 4.31 as compared to 4.35 when using all 7 documents.
 We conclude from these two experiments that LLMs can focus on the beginning and/or the end of their input, but largely ignore the middle.
@@
 4.2 Can we alleviate the middle curse?
-To evaluate loss of performance due to the middle curse in a natural setup, we subsample data points from each of the 5 long-input summarization datasets.
+To evaluate the loss of performance due to the middle curse in a natural setup, we subsample data points from each of the 5 long-input summarization datasets.
 We obtain sentences from the (untruncated) source aligned with the reference summary, following the procedure from Section 3.2.
@@
 We randomly sample 50 data points from each of Arxiv, PubMed, GovReport and Multi-News, and 25 from SummScreenFD, forming an evaluation dataset of 225 samples which we name MiddleSum.
-[^2] We will release it alongside our code
 We evaluate LLMs on MiddleSum, keeping only reference-based evaluation as the dataset is built using saliency with regard to the reference.
@@
 Incremental summarization consists in updating a summary of the text so far with content from the current text block (we have ${\bm{y}}_{0}=\emptyset$):
-$\forall i\in\{1,\ldots,k\},{\bm{y}}_{i}=\text{LLM}({\bm{y}}_{i-1},{\bm{x}}_{i})$ (3)
+${\bm{y}}_{i}$ $=\text{LLM}({\bm{y}}_{i-1},{\bm{x}}_{i})\forall i\in\{1,\ldots,k\}$ (3)
+${\bm{y}}$ $={\bm{y}}_{k}$ (4)
+Figure 6: Long-input summarization performance on Arxiv (top) and GovReport (bottom) with 5 LLMs and all 5 metrics.
+X-axis represents the truncated maximum source length.
+Xgen-7B and Mistral-7B cannot infer beyond 8k tokens.
 In both methods, the final output is ${\bm{y}}$.
@@
 Results are shown in Fig.
-5 (blue and purple bars), with statistics, bigrams analysis and results split over datasets in Appendix E.
-Both methods show promising results, notably on Mistral-7B for which they improve performance significantly.
-Across domains, hierarchical and incremental inference are very successful on scientific publications, which we hypothesize is due to the natural division in structured sections of such inputs.
+5 (blue and purple bars), with detailed numbers in Appendix F.
+We also compare to a baseline consisting in adding the prompt Please also pay attention to the middle section of the input when constructing the summary, which we refer to as the Focus prompt (brown bars).
+Both alternative methods show promising results on open-source LLMs, notably on Mistral-7B for which they improve performance significantly.
+However, they are not successful and lag behind Focus Prompt with GPT-3.5.
+Across domains (see Table 8), hierarchical and incremental inference are very effective on scientific publications, which we hypothesize is due to the natural division in structured sections of such inputs.
 Yet, they seem to harm summaries on the other domains.
 4.3 Is scaling context length really useful?
-Figure 6: Long-input summarization performance on GovReport with 4 models and all 5 metrics.
-X-axis represents the truncated maximum source length.
-Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.
 Experiments from Section 4.1 confirm that LLMs struggle to summarize information contained in the middle of their context window.
-This poses issues for long-input summarization: after the initial part with high saliency, important information becomes sparser, and at the same time LLMs processing capability weakens.
-To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6.
-We use our longest context LLMs Xgen-7B and Mistral-7B, as well as two LLMs extending Llama-2-7B context window with position interpolation (Chen et al., 2023), a method gaining traction as an efficient way to scale LLMs’ context window.
-We use two LLMs applying position interpolation: Vicuna-7B-1.5-16k [^3] In HuggingFace: lmsys/vicuna-7b-v1.5-16k, and Llama-2-7B-32k [^4] In HuggingFace: togethercomputer/LLaMA-2-7B-32K, with context of 16k tokens and 32k tokens, respectively.
-Results for GovReport in Fig.
-6 (and Appendix F for Arxiv and SummScreen) confirm our intuition: all metrics plateau or even decrease (see Mistral-7B) from 4k context window upwards.
-Two conflicting forces are at play when increasing length: giving more information to the model helps it retrieving key elements to make a summary, while reasoning over a longer context is more challenging.
+This poses issues for long-input summarization: after the initial part with (usually) high saliency, important information becomes sparser, and at the same time LLMs processing capability weakens.
+To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6.444Our hardware does not allow us to exceed 12k tokens.
+We use our longest context LLMs {Xgen-7B, Mistral-7B, GPT-3.5} ; as well as two open-source LLMs extending Llama-2-7B context window with position interpolation (Chen et al., 2023), a method gaining traction as an efficient way to scale LLMs’ context window.
+We use Vicuna-7B-1.5-16k [^5] In HuggingFace: lmsys/vicuna-7b-v1.5-16k, and Llama-2-7B-32k [^6] In HuggingFace: togethercomputer/LLaMA-2-7B-32K, with context of 16k tokens and 32k tokens, respectively.
+Results on Arxiv and GovReport in Fig.
+6 confirm our intuition: all metrics plateau or even decrease (see Mistral-7B) from 4k context window upwards.
+Two conflicting forces are at play when increasing length: giving more information to the model helps it retrieve key elements further to make a richer summary, while at the same time reasoning over a longer context is more challenging.
 Yet, such a drop for Xgen-7B and Mistral-7B at 8k inference length is concerning.
-Both position interpolated models show more robustness ; thus these patterns may vary a lot across LLMs.
-Our results suggest that in the current LLMs inference and evaluation framework, there is no need to exceed 4k tokens in the context window.
+Both position interpolated models show more robustness ; while GPT-3.5 seems to plateau at 8k tokens.
+Our results suggest that in the current LLMs inference and evaluation framework, there is no need to exceed 4k tokens in the context window for open-source model.
+4.4 Does the decoding method impact the middle curse?
+We now turn our attention to the process controlling summary generation.
+While we had sampled all summaries with top-k sampling with $k=50$ and temperature $T=0.3$ so far ; we now also experiment with greedy decoding, and top-p sampling (where we use $p=0.95$ and temperature $T=1.0$).
+Figure 7: Summary bigrams (top) and aligned source sentences (bottom) distribution on Arxiv and GovReport for Llama-2-7B and XGen-7B, for several decoding strategies.
+In Fig.
+7, we reproduce the salient bigrams and sentences alignment experiments from Section 3.1 and Section 3.2, respectively, with the aforementioned decoding methods.
+As seen, the decoding method does not affect position bias: for all setups, the LLMs show similar patterns as with our previous default decoding method.
+We conclude that the middle curse is independent from the decoding method.
 5 Related Work
@@
 Liu et al.
-(2023b) proposes to train smaller models like BART (Lewis et al., 2020) or BRIO (Liu et al., 2022b) with contrastive learning using LLMs like ChatGPT as evaluator providing signal on which generated summary candidate is better.
+(2023d) proposes to train smaller models like BART (Lewis et al., 2020) or BRIO (Liu et al., 2022b) with contrastive learning using LLMs like ChatGPT as evaluator providing signal on which generated summary candidate is better.
 Summary chain-of-thought designs a custom chain-of-thought method which first prompts the LLM to list important facts, then integrates these facts into a coherent summary (Wang et al., 2023c).
@@
 It was later found that for in-context learning, the order of examples within the prompt impacts GPT-3’s performance Liu et al.
-(2021); Lu et al.
+(2022a); Lu et al.
 (2021).
@@
 Liu et al.
-(2023a) was the first to show that LLMs’ performance weakens in the middle of the prompt, yet, how LLMs make use of their full context window remains poorly understood.
+(2023a) were the first to show that LLMs’ performance weakens in the middle of the prompt (the middle curse), yet, how LLMs make use of their full context window remains poorly understood.
 The passkey retrieval evaluation, which consists in prompting the LLM to recall a complex string or long number inserted in its prompt, is becoming popular recently as a way of verifying LLM’s processing capability at each position Liu et al.
@@
 6 Conclusion
-Behind the recent hype around LLMs and their amazing instruction following and content generation capacities, our study showcases a major weakness in abstractive summarization: LLMs suffer from the middle curse and struggle to use information in the middle of their context window.
+Behind the recent hype around LLMs and their amazing instruction following and content generation capabilities, our study showcases a major weakness in abstractive summarization: LLMs suffer from the middle curse and struggle to use information in the middle of their context window.
 LLMs do not make a consistent use of their context window as they mostly look at the beginning and (to a lesser extent) the end, which at first glance may be hidden by the prevalent lead bias in summarization datasets.
@@
 We benchmarked two alternative inference methods (hierarchical inference and incremental inference) on MiddleSum, an evaluation subset designed to showcase the middle curse.
-Despite promising results, especially on scientific paper datasets, these methods are far from a silver bullet to the middle curse.
+Despite encouraging results on scientific paper datasets, these methods are far from a silver bullet to the middle curse.
 We call for a better evaluation of LLMs, which accounts for the salient spans of the source which are effectively being processed.
@@
 Findings may change as LLMs undergo changes and improvements in their training process.
+Acknowledgements
+This research was supported by the SINGA scholarship and partially supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.
+Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore.
+We thank anonymous ARR Reviewers 1 and 3 for their great review, which have led us to include plenty more important results enriching our analysis.
+We also thank the AC for their constructive meta-review.
+We thank Florian Le Bronnec and Ruochen Zhao for their helpful proof-reading of the draft.
 Appendix A Statistics
@@
 Underlined test sizes correspond to datasets where we subsample randomly 1,000 test data points for evaluation.
-In Table 4, we include statistics on each of the abstractive summarization datasets under consideration.
-We use the non-anonymized version for CNN/DM (See et al., 2017).
-For Reddit-TIFU, we use the Long subset, and SummScreenFD, is the ForeverDreaming (FD) subset of SummScreen.
-GovReport and SummScreenFD are part of the long-input benchmarks Scrolls (Shaham et al., 2022) and ZeroScrolls (Shaham et al., 2023).
-In Fig. 7, we illustrate how much of the source document(s) is visible with a 4k context window (Llama-2).
-Figure 7: Fraction of the source which fits into the context window, for several context window lengths with Llama-2 tokenization.
-The black dashed lines correspond to Llama-2 context window length of 4k tokens.
-On standard-length datasets, 4k is enough to access 100% of all source documents ; but on the long-input datasets such as GovReport or SummScreenFD, such a context window may not even fit 50% of the source.
-Appendix B GPT-3.5 Evaluation
 Model Metric
@@
 GPT-3.5 3.92 4.30 2.73 3.63 4.47 3.81 2.83 3.63 2.01 1.88 3.60 2.79
-Table 5: Performance achieved by the 5 LLMs on each dataset for 5 automatic metrics.
+[…]
+ROUGE-2 13.17 8.19 5.17 14.47 5.37 9.27 13.78 13.90 17.94 6.55 12.28 12.89
+BERTScore 87.26 87.90 86.38 89.79 83.82 87.03 84.14 84.29 85.99 86.13 86.02 85.31
+A3CU 26.72 15.21 13.37 35.85 7.90 19.81 17.98 18.76 19.32 12.69 22.95 18.34
+SummaC 35.14 23.22 26.00 24.92 37.56 29.37 50.11 47.36 73.23 40.58 48.91 52.04
+GPT-3.5 4.12 4.58 3.08 3.81 4.61 4.04 4.41 4.36 4.03 3.39 4.14 4.07
+Table 5: Performance achieved by the LLMs on each dataset for all 5 metrics.
 # sents is the average number of sentences in generated summaries.
 Multi-X is short for Multi-XScience, Multi-N is Multi-News dataset, AVG columns represent the average over standard-length and long-input datasets, respectively.
-SOTA numbers are taken from(Xie et al., 2023) on CNN/DM, from (Zhao et al., 2022b) XSum, Reddit-TIFU and SAMSum, from (Pang et al., 2022) for Arxiv and PubMed, from (Xiong et al., 2022) for GovReport and SummScreenFD, from (Xiao et al., 2022) for Multi-News and Multi-XScience.
+SOTA numbers are taken from (Xie et al., 2023) on CNN/DM, from (Zhao et al., 2022b) XSum, Reddit-TIFU and SAMSum, from (Pang et al., 2022) for Arxiv and PubMed, from (Xiong et al., 2023) for GovReport and SummScreenFD, from (Xiao et al., 2022) for Multi-News and Multi-XScience.
 Due to a lack of reported results for other metrics, we only include ROUGE-2 scores for SOTA models.
 Best scores (outside of SOTA) are in bold.
+In Table 4, we include statistics on each of the abstractive summarization datasets under consideration.
+We use the non-anonymized version for CNN/DM (See et al., 2017).
+For Reddit-TIFU, we use the Long subset, and SummScreenFD, is the ForeverDreaming (FD) subset of SummScreen.
+GovReport and SummScreenFD are part of the long-input benchmarks Scrolls (Shaham et al., 2022) and ZeroScrolls (Shaham et al., 2023).
+In Fig. 8, we illustrate how much of the source document(s) is visible with a 4k context window (Llama-2).
+Figure 8: Fraction of the source which fits into the context window, for several context window lengths with Llama-2 tokenization.
+The black dashed lines correspond to Llama-2 context window length of 4k tokens.
+On standard-length datasets, 4k is enough to access 100% of all source documents ; but on the long-input datasets such as GovReport or SummScreenFD, such a context window may not even fit 50% of the source.
+Appendix B GPT-3.5 Evaluation
 To evaluate LLM-generated summaries with GPT-3.5, we use the following prompt template:
@@
 Model CNN/DM XSum Reddit-TIFU SAMSum Multi-XScience Arxiv PubMed GovReport SummScreenFD Multi-News
-Flan-UL2 0.000 0.012 0.000 0.017 $0$
+Flan-UL2 0.000 0.012 0.000 0.057 $0$
 Llama-2-7B 0.000 0.000 0.000 0.490 0.000 0.000 0.000 0.000 0.000 0.000
@@
 Mistral-7B 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+GPT-3.5 0.000 0.000 0.000 0.135 0.000 0.000 0.000 0.000 0.000 0.000
 Table 6: P-value of a 2-sample Kolmogorov-Smirnov test between the position distribution of bigrams in LLM-generated summaries and bigrams in reference summaries.
 We round numbers to 3 decimals.
-Numbers in gray correspond to non-significant differences (p-value above 0.05).
+Numbers in gray correspond to non-significant differences (p-value above 0.01).
 Dataset Domain # Docs # Data points # Sentences # Words # Tokens
@@
 Appendix C Baseline Performance
-(a) Llama-2-7B
-(b) Llama-2-13B
-(c) Xgen-7B
-(d) Mistral-7B
-Figure 8: Relative location of summary bigrams within the source on MiddleSum for 4 LLMs.
-In Table 5, we report zero-shot performance with the 5 LLMs described in Section 2.
-We note that for standard-length datasets, Flan-UL2 is dominating, while for long-context ones, Llama-2-13B takes the lead.
+Figure 9: Relative location of summary bigrams within the source on MiddleSum for Llama-2-13B with different inference methods.
+In Table 5, we report zero-shot performance with the 6 LLMs described in Section 2.
+We note that for standard-length datasets, Flan-UL2 is dominating, perhaps due to its better instruction-tuning ; while for long-context ones, GPT-3.5 takes the lead.
+GPT-3.5 is always consistently ahead on the GPT-3.5 metric, showing the preference of the LLM for its own generation when used as an evaluator.
 Appendix D Statistical Significance on RQ1
 In Table 6, we run a 2-sample Kolmogorov-Smirnov statistical significance test to compare the bigrams position distribution of LLMs with the reference summaries.
-Appendix E More Results on MiddleSum
+Appendix E Results with base Models
+Figure 10: Summary bigrams (top) and aligned source sentences (bottom) distribution for Llama-2-7B and Llama-2-13B on the long-input datasets, both with base (dashed lines) and chat (full lines) models.
+In Fig.
+10, we repeat the analysis from Section 3.1 (bigrams alignment) and Section 3.2 (sentences alignment) but this time using the base model versions (not instruction-tuned) of Llama-2-7B and Llama-2-13B.
+As we can see, the position bias is even stronger for base models than for chatbot models, suggesting that this bias is acquired during pre-training.
+Appendix F More Results on MiddleSum
 In Table 7, we show statistics on the MiddleSum evaluation dataset.
-In Figs.
-7(a), 7(b), 7(c) and 7(d), we repeat the salient bigrams analysis from Fig.
-1 for each of the four long-input LLMs on MiddleSum.
-We note that both hierarchical and incremental inference notably decrease reliance on lead bigrams compared to standard inference.
-In Figs. 9, 10, 11, 12 and 13, and in Table 8 (exact numbers), we show reference-based evaluation on each of the 5 subsets of MiddleSum.
-Figure 9: Reference-based evaluation on the Arxiv subset of MiddleSum dataset.
-Figure 10: Reference-based evaluation on the PubMed subset of MiddleSum dataset.
-Figure 11: Reference-based evaluation on the GovReport subset of MiddleSum dataset.
-Figure 12: Reference-based evaluation on the SummScreenFD subset of MiddleSum dataset.
-Figure 13: Reference-based evaluation on the Multi-News subset of MiddleSum dataset.
-Appendix F More Results on Long-Input Tasks
-In Figs. 14 and 15, we conduct the analysis on increasingly longer input context from Fig. 6 on Arxiv and SummScreen, respectively.
-Figure 14: Long-input summarization performance on Arxiv with 4 models and all 5 metrics. X-axis represents the truncated maximum source length. Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.
-Figure 15: Long-input summarization performance on SummScreen with 4 models and all 5 metrics.
-X-axis represents the truncated maximum source length.
-Xgen-7b and Mistral-7b cannot infer beyond 8k tokens.
+In Fig.
+9, we repeat the salient bigrams analysis from Fig.
+1 for Llama-2-13B on MiddleSum.
+We note that both hierarchical and incremental inference notably decrease reliance on lead bigrams compared to standard inference, while the simple Focus Prompt does not.
+In Table 8, we show reference-based evaluation on each of the 5 subsets of MiddleSum.
 Model Metric Inference MiddleSum (MS) MS/Arxiv MS/PubMed MS/GovReport MS/SummScreen MS/Multi-News
 Llama-2-7B ROUGE-2 Standard 10.96 12.62 10.97 13.26 4.07 10.43
+Focus prompt 10.70 12.10 11.51 13.28 3.57 9.48
 Hierarchical 11.33 14.63 13.36 13.51 4.85 7.06
@@
 […]
+Focus prompt 84.07 83.33 83.00 84.77 85.27 84.59
 Hierarchical 84.13 83.86 83.74 84.26 85.23 84.10
@@
 […]
+Focus prompt 12.41 13.54 13.39 10.84 7.15 14.52
 Hierarchical 13.21 15.64 15.14 12.26 9.71 11.57
@@
 Llama-2-13B ROUGE-2 Standard 11.07 11.68 11.63 13.56 5.09 10.38
+Focus prompt 10.82 11.51 11.30 12.42 4.83 11.03
 Hierarchical 10.24 13.45 13.26 10.21 4.93 6.71
@@
 […]
+Focus prompt 84.12 83.26 83.06 84.49 85.08 85.19
 Hierarchical 83.17 83.74 83.75 80.07 85.03 84.20
@@
 […]
+Focus prompt 13.10 13.17 11.76 11.91 7.06 18.58
 Hierarchical 12.85 16.28 14.11 10.40 8.36 12.86
@@
 Xgen-7B ROUGE-2 Standard 8.92 11.94 9.34 10.02 4.53 6.57
+Focus prompt 9.79 12.87 10.79 11.46 4.15 6.85
 Hierarchical 9.87 13.06 11.37 11.77 2.79 6.83
@@
 […]
+Focus prompt 82.75 82.45 82.14 83.08 83.81 82.80
 Hierarchical 82.86 82.73 82.56 82.67 83.79 83.03
@@
 […]
+Focus prompt 12.23 15.82 12.04 11.57 8.20 11.50
 Hierarchical 11.74 14.55 12.61 10.76 7.33 11.25
@@
 Mistral-7B ROUGE-2 Standard 7.16 9.28 9.56 5.86 2.19 6.41
+Focus prompt 6.20 7.62 8.44 4.85 1.42 6.29
 Hierarchical 10.78 11.67 13.71 13.36 4.76 7.39
@@
 […]
+Focus prompt 80.72 80.31 80.81 79.91 80.12 82.13
 Hierarchical 83.10 82.56 83.00 83.81 83.90 82.62
@@
 […]
+Focus prompt 9.96 10.13 10.65 7.81 7.59 12.44
 Hierarchical 11.15 12.15 14.21 10.62 7.48 9.47
 Incremental 10.43 14.25 11.00 8.90 7.44 9.05
+GPT-3.5 ROUGE-2 Standard 12.33 12.99 12.78 16.16 6.05 10.53
+Focus prompt 12.17 12.61 12.56 15.95 6.35 10.46
+Hierarchical 8.59 9.58 10.27 10.30 4.33 6.36
+Incremental 12.04 11.85 12.17 18.46 5.38 9.03
+[…]
+Focus prompt 84.80 83.76 83.81 85.50 86.28 85.39
+Hierarchical 84.27 83.34 83.62 85.30 85.37 84.28
+Incremental 84.43 83.48 83.74 85.41 85.76 84.44
+[…]
+Focus prompt 15.07 14.51 15.59 12.34 13.63 18.54
+Hierarchical 12.24 12.54 13.01 13.29 9.03 11.74
+Incremental 12.08 13.59 13.24 10.90 9.05 12.08
 Table 8: Reference-based results for all models and inference methods on MiddleSum, breaking down by subset. The best number across inference methods is in bold.

### v6

@@
 This setup enables us to shuffle the order of the input, which is not realistic for the single-document setup as it would break coherence.
-We only consider data points with the same number $k$ of documents: $k=7$ documents for Multi-XScience ($n=329$), and $k=5$ documents for Multi-News ($n=219$).
+We only consider data points with the same number $k$ of documents: $k=7$ documents for Multi-XScience ($n=329$), and $k=5$ documents for Multi-News ($n=219$) [^4] We don’t subsample from these subsets for GPT-3.5..
 In the first experiment, we vary the position of salient information throughout the input.
@@
 R-2 is ROUGE-2, BS refers to BERTScore.
-% is the mean relative performance compared to the baseline with all documents.
+% is the mean relative performance across all metrics compared to the baseline with all documents.
 Figure 5: Reference-based evaluation on the MiddleSum dataset.
@@
 This poses issues for long-input summarization: after the initial part with (usually) high saliency, important information becomes sparser, and at the same time LLMs processing capability weakens.
-To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6.444Our hardware does not allow us to exceed 12k tokens.
+To investigate this issue, we infer long-document summarization with length truncated at $m*2k$ tokens, varying $m$ from 1 to 6.555Our hardware does not allow us to exceed 12k tokens.
 We use our longest context LLMs {Xgen-7B, Mistral-7B, GPT-3.5} ; as well as two open-source LLMs extending Llama-2-7B context window with position interpolation (Chen et al., 2023), a method gaining traction as an efficient way to scale LLMs’ context window.
@@
 Limitations
-Our work only considers open-source LLMs for summary generation and ignores closed-source LLMs such as the popular OpenAI’s GPT-3.5 and GPT-4, or Anthropic’s Claude.
-We made this decision to advocate for openness in LLM research ; yet we acknowledge that it would be interesting to also investigate properties of summaries generated by these paying API LLMs.
+Our work mainly considers open-source LLMs for summary generation and ignores some popular closed-source LLMs such as Anthropic’s Claude.
+We made this decision to advocate for openness in LLM research ; yet we acknowledge that it would be interesting to also investigate properties of summaries generated by more paying API LLMs.
 Another limitation lays in the saliency estimation.
@@
 Other methods are also well-suited for this task, albeit at greater computational cost ; for instance semantic similarity through BERTScore or BARTScore ; or saliency estimation through a LLM in zero-shot.
-Lastly, we can only evaluate a finite number of LLMs, and we settled for the evaluation of 5 recent and popular open-source LLMs.
-Findings may change as LLMs undergo changes and improvements in their training process.
+Lastly, we can only evaluate a finite number of LLMs, and we settled for the evaluation of 6 recent and popular LLMs.
+Findings may change as LLMs undergo changes and improvements in their pre-training and instruction-tuning methodologies.
 Acknowledgements
@@
 Reddit-TIFU (Long) (Kim et al., 2019) Social Media 1.00 33,701 4,214 4,221 22.21 1.45 2 444.20 23.37 532.18 29.82 128
-SAMSum Kim et al. (2019) Dialogue 1.00 14,732 818 819 8.96 2.01 2 126.93 23.12 175.54 29.69 128
+SAMSum Gliwa et al. (2019) Dialogue 1.00 14,732 818 819 8.96 2.01 2 126.93 23.12 175.54 29.69 128
 Multi-XScience (Lu et al., 2020) Science 5.06 30,369 5,066 5,093 30.55 4.86 5 773.36 120.65 965.99 157.77 384